OpenStudy (anonymous):

26, 10, 17 right, acute, or obtuse.

OpenStudy (anonymous):

26, 14, 22 right, acute, or obtuse. Is the nest problem if you could please :)

OpenStudy (campbell_st):

use pythagoras' theorem find 26^2 = 10^2 + 17^2 then use the conditions below... c^2> a^2 + b^2 the triangle is obtuse angled c^2 < a^2 + b^2 the triangle is acute angled c^2 = a^2 + b^2 the triangle is right angled.

OpenStudy (anonymous):

wow, what did you just do there?

OpenStudy (omnomnom):

He helped you.. DUH!

OpenStudy (campbell_st):

ok... what is 26^2? and what is 10^2 + 17^2 ?

OpenStudy (campbell_st):

you need to let c = 26, a = 10 and b = 17